the perimeter of a rectangular lot is 190 ft. the width of the lot is 15ft more than the length. what are the dimensions of the lot.
P=2L+2W
W=L+15
190=2L+2(L+15)
190=2L+2L+30
190-30=4L
160/4=L
40=L
W=40+15=55
DIMENSIONS IS 55 FEET BY 40 FEET.

Answer by chiexpert(48)

P=2L+2W
W=L+15
190=2L+2(L+15)
190=2L+2L+30
190-30=4L
160/4=L
40=L
W=40+15=55
DIMENSIONS IS 55 FEET BY 40 FEET.